calculate aqueous solubility of hexachlorobenzene in water using
a) molecular area 203 angstrom
b) molecular volume 171 angstrom
c) molar volume 181 cm3/mol
Answer #1

(a). Molecular area of hexachlorobenzene = 203 Ao2

For hexachlorobenzene -

a1 = 3.29
a2 = - 0.0422
a3 = - 0.0103

Tm for pure hexachlorobenzene = 231 oC

log C = a1 + a2*Am + a3*Tm

= 3.29 - 0.0422 * 203 - 0.0103 * 231

= 3.29 - 8.566 - 2.3793

= - 7.6553

C = 10^-7.6553

= 2.211 x 10-8 mol/L